# A Variational Framework for the Algorithmic Complexity of PDE Solutions

### v1

@@ 1 Introduction
For centuries, the modeling and description of physical phenomena have relied on the formulation and analysis of Partial Differential Equations (PDEs). While classical analytical methods have provided closed-form solutions for a limited number of cases, the vast majority of physically relevant PDEs require numerical methods to approximate their solutions and simulate the underlying systems. Entire areas within scientific computing and numerical analysis focus on designing and analyzing these numerical methods, ensuring the well-posedness of the problem, and the accuracy and efficiency of the resulting algorithms. Yet, the fundamental question of Turing computability [1] of PDE solutions has only been addressed to a limited extent. In this context, the computability question is, however, relevant because digital computers are the only hardware platforms for numerical calculations, with Turing machines defining their theoretical limits. Thus, tasks that are intractable for Turing machines are necessarily more challenging for digital computers in practice. The central problem is to determine whether the infinite-dimensional solutions of PDEs can, in principle, be approximated effectively on a digital machine and, if so, what computational resources are inherently required. Within the framework of computability analysis and complexity theory, we can distinguish PDEs that are effectively solvable from those that may encode undecidable behavior and quantify the efficiency of approximation when computability is assured. Therefore, addressing the computability and complexity of PDE solutions not only clarifies the limits of current digital hardware for solving PDEs but also highlights classes of equations where novel computational architectures [2] and new notions of computability may be necessary to ensure feasible algorithms for solving such PDEs.
@@
Classical analyzes of computability and complexity of PDEs have primarily addressed equations admitting closed-form solutions [3, 4, 5]. However, for most PDEs arising in real-world applications, such explicit solutions are unavailable [6, 7], motivating the development of new frameworks to study their computability and complexity. Recent advances in variational methods, especially least-squares formulations [8, 9], offer a flexible framework that has already shown success in nonlinear, high-dimensional, and even discontinuous problems [10, 11, 12].
@@
In this work, we leverage the gradient flow associated with least-squares variational formulations [13, 14] to develop a novel framework for analyzing the computability and complexity of PDE solutions. Our approach establishes a direct link between the complexity of solving a PDE and the convergence rates of its discrete gradient flow. This connection relates structural properties of the PDE operator—such as coercivity, ellipticity, and convexity—to the complexity class of its solutions, thereby extending existing results to a much broader class of PDEs. Moreover, it establishes a relation between the loss of regularity of the PDE solutions with the complexity blowup, where polynomial-time input data produce solutions with super-polynomial complexity [5].
@@
This framework contributes to the development of a more general computability and complexity theory for PDEs, extending to both linear and nonlinear equations without closed-form solutions. Its relevance is illustrated through two case studies: the polynomial time computability of the Poisson equation [6] and the intrinsic complexity blowup of the Eikonal equation [15].
@@ 2 Related Work
The literature on complexity theory for PDEs focuses mainly on four complexity classes [5]. The class $P$ consists of all decision problems that can be solved in polynomial time by a deterministic Turing machine. The class $NP$ contains those problems for which a proposed solution can be verified in polynomial time. The class #P\#P encompasses counting problems, namely, those for which the number of solutions to an $NP$ problem can be computed. Finally, the class $FP$ includes all the counting problems that can be solved by a function-oracle Turing machine in polynomial time [16].
@@
The arguably first study on the computability of PDE solutions, presented in [17], examined a wave equation of the form
@@
Additional work in [19] addressed the inhomogeneous linear Schrödinger equation
@@
showing that the solution $u\in H^{s}(\mathbb{R}^{d}\times(0,\infty))$ is computable for all $\phi\in C^{0}(\mathbb{R}^{d}\times(0,\infty))$. Moreover, in the homogeneous case ($\phi=0$), it was shown that the solution $u\in L^{p}(\mathbb{R}^{d}\times(0,\infty))$ is computable if and only if $p=2$. Extensions to the nonlinear setting are presented in [4], establishing the computability of solutions in $H^{1}(\mathbb{R}^{d}\times(0,\infty))$.
@@
It was shown in [20] that if the complexity classes $FP$ and #P\#P coincide (a stronger assumption than $P=NP$), then any solution $u\in C^{2}(B_{1})$ over the unit ball $B_{1}:=\{x\in\mathbb{R}^{d}:|x|_{2}\leq 1\}$ is polynomial time computable, as formalized in Definition 6.4. This holds whenever $f$ and $g$ are smooth and polynomial time computable on $B_{1}$ and $\partial B_{1}$, respectively.
The results for the Laplace and diffusion equations were recently improved in [5], showing that the solution in each case is #P\#P-complete when the initial and boundary data are polynomial time computable. Moreover, it was shown in [5] that these equations exhibit a complexity blowup, implying that no numerical algorithm can compute their solutions in polynomial time for arbitrary initial and boundary data, unless FP=#PFP=\#P.
@@
with $\mathcal{A}$ is a linear differential operator, $\phi$ the initial condition and $\mathcal{L}$ is a linear operator describing the boundary conditions. In [21], it is shown that the solution $u$ of (8) is computable in $PSPACE$ [22] under certain conditions on the initial and boundary data. In particular, the differential operator $\mathcal{A}$ must be computable in polynomial time, with the solution admitting a finite-difference approximation $u_{n}$ converging to the solution $u$ in the supremum norm.
The contributions of the present paper follow a similar philosophy to [21], where we estimate the complexity of a numerical algorithm approximating the solution of the PDE to characterize the intrinsic complexity of the solution itself. Our approach, however, introduces a novel and more general framework based on discrete gradient flows, offering a powerful characterization of the complexity of broad classes of PDEs and identifying regimes in which their solutions exhibit complexity blowup.
@@ 3 Contributions
In this paper, we present a novel variational framework for analyzing the complexity of different classes of PDEs by considering a variational formulation whose minimizer solves the PDE, and estimating the complexity of its discrete gradient flow. The theoretical framework developed here builds upon previous work on Sobolev cubatures [11] and polynomial surrogate models for PDE learning problems [23]. The main contributions on this paper are summarized as follows:
@@ 4 Variational Framework
The proof of the error inequality in (23) follows directly from the quadratic growth condition together with the triangle inequality; see [12] for further details. The exponential convergence of the Euler gradient flow is then obtained by applying Lemma 4.11 in combination with the classic result for Lipschitz-smooth functionals satisfying the RSI condition [24]. ∎
@@
It is a standard result [25] that the exponential decay of the optimization error in (25) is guaranteed whenever the functional $\mathcal{L}:\mathcal{S}_{\theta}(\Omega)\to\mathbb{R}_{+}$ is $\lambda$-convex, since $\lambda$-convexity is a stronger condition than the QGC.
@@ 5 Finite-Dimensional Approximations
A proof of this result can be found in the 1-dimensional setting in [26] and in higher dimensions in [27]. ∎
@@
Let $E_{\rho_{i}}\in\mathbb{C}$ be the $i$-th Bernstein Ellipse [28], defined as the ellipse in the complex $x_{i}$-plane with foci 1-1 and $1$ and radius $\rho_{i}\in\mathbb{R}_{+}$ for $i\in\{1,\dots,d\}$. The elliptic polycylinder $E(\rho)\subseteq\mathbb{C}^{d}$, $\rho\in\mathbb{R}^{d}$, is defined as
@@
[29] Let $f\in AC^{k-1}(\Omega)\cap BV^{k}(\Omega)$ be a function with $k$-th order bounded variations and k1k-1 absolutely continuous derivatives, then the Chebyshev approximation $f_{n}\in\Pi_{n}(\Omega)$ converges with the rate
@@
The proof of Propositions 5.5 and  5.6 can be found in 1-D in [28] and in general dimensions in [12] and [29]. ∎
In Section 6 we will see how Proposition 5.6 is not sufficient for showing the complexity blowup of PDEs, as it is only an upper bound on the convergence rate. Therefore, following [26, 15], we recall the following reverse theorem.
Theorem 5.7 ([26]).
@@
Let $P_{n}(\Omega)\subseteq\Omega$ be the Legendre points [30] re-scaled to $\Omega$. The Lagrange polynomials [26] $\{L_{\alpha}\}_{\alpha\in A_{n,d}}\subseteq\Pi_{n}(\Omega)$ defined with respect to the Legendre grid, are given by
@@
To approximate the PDE learning losses in (11), we introduce the Sobolev cubatures following [23, 11].
@@
[12] Let $f\in H^{k}(\Omega)$ be a Sobolev function. If in addition $f\in AC^{k-1}(\Omega)$, then the $s$-th order Sobolev cubature of degree $n\in\mathbb{N}$ converges to the squared Sobolev norm $I^{s}[f]:=\|f\|_{H^{s}(\Omega)}^{2}$, for $0\leq s\leq k$, with the rate
@@
where $V_{k}:=\mathcal{V}_{k}(f)$ are the $k$-th order variations of $f$ [23].
@@
The proof of this result can be found in [23]. ∎
@@ 6 Computability Theory in Banach Spaces
In this section, we introduce the notion of Turing computability in Banach spaces, following [31], and discuss some of its analytical properties. Our focus centers on computability in the space of continuous functions $C^{0}(\Omega)$ and the Sobolev spaces $H^{k}(\Omega)$ of order $k\in\mathbb{R}$. Let $\Omega:=(-1,1)^{d}\subseteq\mathbb{R}^{d}$ be the computable hypercube of dimension $d$, We begin by revisiting the concept of $C^{0}$-computability.
@@
We emphasize that, as shown in [31], Definition 6.1 is equivalent to the classical notion of Turing-computable functions [32], namely functions that are sequentially computable and effectively uniformly continuous. We omit other characterizations of $C^{0}$-computability, since the effective Weierstrass formulation in Definition 6.1 is naturally aligned with the approximation-theoretic perspective of our framework.
@@
Using the standard result that $C^{0}_{c}(\Omega)$ spaces are dense in $L^{2}(\Omega)$ [33], it follows that Sobolev spaces are a natural example of Banach spaces equipped with a computability structure.
@@ 6.1 Complexity of PDE Solutions
We start by noting that since the functional $\mathcal{L}:H^{k}(\Omega)\rightarrow\mathbb{R}_{+}$ is lower semi-continuous, convex, and coercive, hence the existence of a minimizer $u^{*}\in H^{k}(\Omega)$ is guaranteed by applying Tonelli’s direct method [34].
@@
For the second part of the statement, we apply Lemma 6.15, which ensures that the Dirichlet boundary condition $g$ and the source term $f$ are computable in polynomial time. Furthermore, by Theorem 6.13 the solution is $H^{k}$-computable. However, by Theorem 5.7 together with the Sobolev embedding theorem [35] for $k>d/2$, there exists a sub-exponential function $\mathcal{Q}:\mathbb{N}\to\mathbb{R}+$ such that
@@ 7.1 Complexity of solutions to the Poisson Equation
Next, we state the result, ensuring the well-posedness of the Poisson learning problem. A similar assertion can be found in [36].
@@
Note that, in general, the assumptions in Theorem 7.3 are not sufficient to ensure that the solution $u^{*}$ inherits the regularity of the data, due to the presence of corners in the hypercube $1,1]^{d}$. To address this issue, we assume additional compatibility conditions [37] on the boundary term $g$. A detailed analysis of how the regularity of the domain influences the complexity of the solution will be the subject of future work.
@@
as shown in [38] and $\|u\|_{H^{-1}(\Omega)}^{2}\leq\|u\|_{L^{2}(\Omega)}^{2}$ by definition of the negative order norm, we obtain
@@
To finish the proof, we use that the Laplace operator is a hypo-elliptic operator [39], hence, the solution $u^{*}\in H^{1}(\Omega)$ inherits the regularity of the domain, boundary condition and right hand-side. Therefore, as a result of Corollary 6.16, we can characterize the complexity of the solution as follows:

### v2

@@ 1 Introduction
For centuries, the modeling and description of physical phenomena have relied on the formulation and analysis of Partial Differential Equations (PDEs). While classical analytical methods have provided closed-form solutions for a limited number of cases, the vast majority of physically relevant PDEs require numerical methods to approximate their solutions and simulate the underlying systems. Entire areas within scientific computing and numerical analysis focus on designing and analyzing these numerical methods, ensuring the well-posedness of the problem, and the accuracy and efficiency of the resulting algorithms. Yet, the fundamental question of Turing computability [Church1937-sr] of PDE solutions has only been addressed to a limited extent. In this context, the computability question is, however, relevant because digital computers are the only hardware platforms for numerical calculations, with Turing machines defining their theoretical limits. Thus, tasks that are intractable for Turing machines are necessarily more challenging for digital computers in practice. The central problem is to determine whether the infinite-dimensional solutions of PDEs can, in principle, be approximated effectively on a digital machine and, if so, what computational resources are inherently required. Within the framework of computability analysis and complexity theory, we can distinguish PDEs that are effectively solvable from those that may encode undecidable behavior and quantify the efficiency of approximation when computability is assured. Therefore, addressing the computability and complexity of PDE solutions not only clarifies the limits of current digital hardware for solving PDEs but also highlights classes of equations where novel computational architectures [Handler1967-ty] and new notions of computability may be necessary to ensure feasible algorithms for solving such PDEs.
@@
Classical analyzes of computability and complexity of PDEs have primarily addressed equations admitting closed-form solutions [Pour-El1997-br, Weihrauch2006-cp, bacho2023complexityblowupsolutionslaplace]. However, for most PDEs arising in real-world applications, such explicit solutions are unavailable [Evans2022-ig, Trefethen2000-su], motivating the development of new frameworks to study their computability and complexity. Recent advances in variational methods, especially least-squares formulations [Los2023-jo, Cai2020-js], offer a flexible framework that has already shown success in nonlinear, high-dimensional, and even discontinuous problems [Hu2024-sy, Cardona2023-ma, Suarez2025].
@@
In this work, we leverage the gradient flow associated with least-squares variational formulations [Ghoussoub2004-ou, Ambrosio2008-qd] to develop a novel framework for analyzing the computability and complexity of PDE solutions. Our approach establishes a direct link between the complexity of solving a PDE and the convergence rates of its discrete gradient flow. This connection relates structural properties of the PDE operator—such as coercivity, ellipticity, and convexity—to the complexity class of its solutions, thereby extending existing results to a much broader class of PDEs. Moreover, it establishes a relation between the loss of regularity of the PDE solutions with the complexity blowup, where polynomial-time input data produce solutions with super-polynomial complexity [bacho2023complexityblowupsolutionslaplace].
@@
This framework contributes to the development of a more general computability and complexity theory for PDEs, extending to both linear and nonlinear equations without closed-form solutions. Its relevance is illustrated through two case studies: the polynomial time computability of the Poisson equation [Evans2022-ig] and the intrinsic complexity blowup of the Eikonal equation [Alimov2019-np].
@@ 2 Related Work
The literature on complexity theory for PDEs focuses mainly on four complexity classes [bacho2023complexityblowupsolutionslaplace]. The class $P$ consists of all decision problems that can be solved in polynomial time by a deterministic Turing machine. The class $NP$ contains those problems for which a proposed solution can be verified in polynomial time. The class #P\#P encompasses counting problems, namely, those for which the number of solutions to an $NP$ problem can be computed. Finally, the class $FP$ includes all the counting problems that can be solved by a function-oracle Turing machine in polynomial time [Arora2012-bh].
@@
The arguably first study on the computability of PDE solutions, presented in [Pour-El1983-ls], examined a wave equation of the form
@@
Additional work in [Weihrauch2001-rm] addressed the inhomogeneous linear Schrödinger equation
@@
showing that the solution $u\in H^{s}(\mathbb{R}^{d}\times(0,\infty))$ is computable for all $\phi\in C^{0}(\mathbb{R}^{d}\times(0,\infty))$. Moreover, in the homogeneous case ($\phi=0$), it was shown that the solution $u\in L^{p}(\mathbb{R}^{d}\times(0,\infty))$ is computable if and only if $p=2$. Extensions to the nonlinear setting are presented in [Weihrauch2006-cp], establishing the computability of solutions in $H^{1}(\mathbb{R}^{d}\times(0,\infty))$.
@@
It was shown in [Kawamura2017-dv] that if the complexity classes $FP$ and #P\#P coincide (a stronger assumption than $P=NP$), then any solution $u\in C^{2}(B_{1})$ over the unit ball $B_{1}:=\{x\in\mathbb{R}^{d}:|x|_{2}\leq 1\}$ is polynomial time computable, as formalized in Definition 6.4. This holds whenever $f$ and $g$ are smooth and polynomial time computable on $B_{1}$ and $\partial B_{1}$, respectively.
The results for the Laplace and diffusion equations were recently improved in [bacho2023complexityblowupsolutionslaplace], showing that the solution in each case is #P\#P-complete when the initial and boundary data are polynomial time computable. Moreover, it was shown in [bacho2023complexityblowupsolutionslaplace] that these equations exhibit a complexity blowup, implying that no numerical algorithm can compute their solutions in polynomial time for arbitrary initial and boundary data, unless FP=#PFP=\#P.
@@
with $\mathcal{A}$ is a linear differential operator, $\phi$ the initial condition and $\mathcal{L}$ is a linear operator describing the boundary conditions. In [Koswara2021-sr], it is shown that the solution $u$ of (8) is computable in $PSPACE$ [Arora2009-ji] under certain conditions on the initial and boundary data. In particular, the differential operator $\mathcal{A}$ must be computable in polynomial time, with the solution admitting a finite-difference approximation $u_{n}$ converging to the solution $u$ in the supremum norm.
The contributions of the present paper follow a similar philosophy to [Koswara2021-sr], where we estimate the complexity of a numerical algorithm approximating the solution of the PDE to characterize the intrinsic complexity of the solution itself. Our approach, however, introduces a novel and more general framework based on discrete gradient flows, offering a powerful characterization of the complexity of broad classes of PDEs and identifying regimes in which their solutions exhibit complexity blowup.
@@ 3 Contributions
In this paper, we present a novel variational framework for analyzing the complexity of different classes of PDEs by considering a variational formulation whose minimizer solves the PDE, and estimating the complexity of its discrete gradient flow. The theoretical framework developed here builds upon previous work on Sobolev cubatures [Cardona2023-ma] and polynomial surrogate models for PDE learning problems [Cardona2024-sr]. The main contributions on this paper are summarized as follows:
@@ 4 Variational Framework
The proof of the error inequality in (23) follows directly from the quadratic growth condition together with the triangle inequality; see [Suarez2025] for further details. The exponential convergence of the Euler gradient flow is then obtained by applying Lemma 4.11 in combination with the classic result for Lipschitz-smooth functionals satisfying the RSI condition [Karimi2016-zl]. ∎
@@
It is a standard result [karimi2020] that the exponential decay of the optimization error in (25) is guaranteed whenever the functional $\mathcal{L}:\mathcal{S}_{\theta}(\Omega)\to\mathbb{R}_{+}$ is $\lambda$-convex, since $\lambda$-convexity is a stronger condition than the QGC.
@@ 5 Finite-Dimensional Approximations
A proof of this result can be found in the 1-dimensional setting in [Trefethen2019] and in higher dimensions in [MASON1980349]. ∎
@@
Let $E_{\rho_{i}}\in\mathbb{C}$ be the $i$-th Bernstein Ellipse [Trefethen2017_mD], defined as the ellipse in the complex $x_{i}$-plane with foci 1-1 and $1$ and radius $\rho_{i}\in\mathbb{R}_{+}$ for $i\in\{1,\dots,d\}$. The elliptic polycylinder $E(\rho)\subseteq\mathbb{C}^{d}$, $\rho\in\mathbb{R}^{d}$, is defined as
@@
[zavalani2024] Let $f\in AC^{k-1}(\Omega)\cap BV^{k}(\Omega)$ be a function with $k$-th order bounded variations and k1k-1 absolutely continuous derivatives, then the Chebyshev approximation $f_{n}\in\Pi_{n}(\Omega)$ converges with the rate
@@
The proof of Propositions 5.5 and  5.6 can be found in 1-D in [Trefethen2017_mD] and in general dimensions in [Suarez2025] and [zavalani2024]. ∎
In Section 6 we will see how Proposition 5.6 is not sufficient for showing the complexity blowup of PDEs, as it is only an upper bound on the convergence rate. Therefore, following [Trefethen2019, Alimov2019-np], we recall the following reverse theorem.
Theorem 5.7 ([Trefethen2019]).
@@
Let $P_{n}(\Omega)\subseteq\Omega$ be the Legendre points [Golub1969-sb] re-scaled to $\Omega$. The Lagrange polynomials [Trefethen2019] $\{L_{\alpha}\}_{\alpha\in A_{n,d}}\subseteq\Pi_{n}(\Omega)$ defined with respect to the Legendre grid, are given by
@@
To approximate the PDE learning losses in (11), we introduce the Sobolev cubatures following [Cardona2024-sr, Cardona2023-ma].
@@
[Suarez2025] Let $f\in H^{k}(\Omega)$ be a Sobolev function. If in addition $f\in AC^{k-1}(\Omega)$, then the $s$-th order Sobolev cubature of degree $n\in\mathbb{N}$ converges to the squared Sobolev norm $I^{s}[f]:=\|f\|_{H^{s}(\Omega)}^{2}$, for $0\leq s\leq k$, with the rate
@@
where $V_{k}:=\mathcal{V}_{k}(f)$ are the $k$-th order variations of $f$ [Cardona2024-sr].
@@
The proof of this result can be found in [Cardona2024-sr]. ∎
@@ 6 Computability Theory in Banach Spaces
In this section, we introduce the notion of Turing computability in Banach spaces, following [Pour-El1989-yg], and discuss some of its analytical properties. Our focus centers on computability in the space of continuous functions $C^{0}(\Omega)$ and the Sobolev spaces $H^{k}(\Omega)$ of order $k\in\mathbb{R}$. Let $\Omega:=(-1,1)^{d}\subseteq\mathbb{R}^{d}$ be the computable hypercube of dimension $d$, We begin by revisiting the concept of $C^{0}$-computability.
@@
We emphasize that, as shown in [Pour-El1989-yg], Definition 6.1 is equivalent to the classical notion of Turing-computable functions [Grzegorczyk1957], namely functions that are sequentially computable and effectively uniformly continuous. We omit other characterizations of $C^{0}$-computability, since the effective Weierstrass formulation in Definition 6.1 is naturally aligned with the approximation-theoretic perspective of our framework.
@@
Using the standard result that $C^{0}_{c}(\Omega)$ spaces are dense in $L^{2}(\Omega)$ [Rudin1986-qe], it follows that Sobolev spaces are a natural example of Banach spaces equipped with a computability structure.
@@ 6.1 Complexity of PDE Solutions
We start by noting that since the functional $\mathcal{L}:H^{k}(\Omega)\rightarrow\mathbb{R}_{+}$ is lower semi-continuous, convex, and coercive, hence the existence of a minimizer $u^{*}\in H^{k}(\Omega)$ is guaranteed by applying Tonelli’s direct method [Maso1993-ss].
@@
For the second part of the statement, we apply Lemma 6.15, which ensures that the Dirichlet boundary condition $g$ and the source term $f$ are computable in polynomial time. Furthermore, by Theorem 6.13 the solution is $H^{k}$-computable. However, by Theorem 5.7 together with the Sobolev embedding theorem [brezis2011] for $k>d/2$, there exists a sub-exponential function $\mathcal{Q}:\mathbb{N}\to\mathbb{R}+$ such that
@@ 7.1 Complexity of solutions to the Poisson Equation
Next, we state the result, ensuring the well-posedness of the Poisson learning problem. A similar assertion can be found in [Bochev2009-ut].
@@
Note that, in general, the assumptions in Theorem 7.3 are not sufficient to ensure that the solution $u^{*}$ inherits the regularity of the data, due to the presence of corners in the hypercube $1,1]^{d}$. To address this issue, we assume additional compatibility conditions [Hell2014-wc] on the boundary term $g$. A detailed analysis of how the regularity of the domain influences the complexity of the solution will be the subject of future work.
@@
as shown in [Diethelm2002-kr] and $\|u\|_{H^{-1}(\Omega)}^{2}\leq\|u\|_{L^{2}(\Omega)}^{2}$ by definition of the negative order norm, we obtain
@@
To finish the proof, we use that the Laplace operator is a hypo-elliptic operator [Mitrea2013-cl], hence, the solution $u^{*}\in H^{1}(\Omega)$ inherits the regularity of the domain, boundary condition and right hand-side. Therefore, as a result of Corollary 6.16, we can characterize the complexity of the solution as follows: